From GlueXWiki
Jump to: navigation, search

September 13, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
  2. Cathode corrosion
    • Recent tests
  3. Engineering (Bill)
  4. Electronics (Chris, Nick)
  5. Third package tests FDC E-log (Beni)
  6. TDR, pages 132-143 (Lubomir)
  7. Other


Participants: Eugene, Bill, Dave, Simon, Chris, Vlad, Beni, and Lubomir


- Dave: 7 cathodes ready, another 5 need daughter cards to be glued. Will have 200 new cards (Lubomir) beginning of next month, we better use them for the rest of the cathodes; so far we have 3 cathodes with refurbished daughter cards. One new wire frame (#29) is ready for stringing, Anatoly just put HV caps and tested it. Adam (Vision Machine) will bring one g10 wire frame next Wednesday and another three on Friday. After Rohacell lamination Adam will fly cut them and we will have the first frame for the PCB lamination next Friday. Bill: Adam will need our de-humidifier.

- Lubomir: we may try to fix the wire frames #27 and #28 that are drawing high current. Suspect the via of the HV trace that connects the bottom trace just against the Rohacell. If this is the problem, can be fixed by drilling a hole from the back side. Bill agreed and we will try it next month when we will have time.

Cathode corrosion

- On Monday we started a test with a cathode (the demo one) using two techniques to protect the copper from corrosion: one side lightly coated with Hysol, the other with Kapton films sitting on Apiezon. We installed this cathode in a cell with Viton+Apiezon and started flushing with gas at 200ccpm (to have enough flow through the oxygen sensor). Some leakage was found (< 10^-4) at the Kapton side but that pressure inside is very high due to the higher flow rate per cell. Interestingly leakages were found at the two transition regions between the Kapton and Hysol areas, probably the Hysol layer is thicker than the Kapton film. Importantly, the oxygen went down to 180ppm in three days.

- We will wait till Monday to see the final oxygen percentage and will open the test package. If everything is OK we want to use the Hysol technique for the first half of the first package. Bill wanted to know if Dave is happy with the Hysol coating, then they discussed how to improve the flatness of the Hysol layer. In any case we will practice more before working on the production cathodes. Bill also proposed to use conformal coating instead of Hysol, but Eugene and Lubomir didn't agree since it introduces new material and there are complains from drift chamber groups about possible out-gassing. Even we have some parts of the PCBs coated inside the gas volume and we try to find the same material, we still need to do long tests before we are sure it will not harm the chamber. As for the other technique, using Apiezon on big areas it's messy and not possible to clean, it may start leaking at higher temperature when chamber is in vertical position?

- Vlad started a new test in the vacuum chamber now adding a sample with Viton+Apiezon against cathode foil that was coated with Hysol. The chamber will stay at very high temperature (close to 100deg C) and 100% humidity till Monday.


- Bill answering Eugene's questions: the three carts that will support FDC, CDC and the FDC cables will be ready in a couple of weeks. The fiducials will be glued to the gusset ring and then surveyed w.r.t. wire planes and the cart at Blue Crab. Another survey will be done when the cart is at the platform and another one of the optical targets at the front and back of the FDC when the detector is installed. Bill will estimate the man-weeks for the surveys.

- When there's time in the next two months, Dave will start working on the cooling system of the third package in 126. Bill will make sure we have all the parts and tools.


- Nick: by next Wednesday we will have another set of PCBs populated. We will start testing the LV system, Beni needs two LV cables for the two fADC125 he's planing to use in the DAQ.

Third package tests

- Beni and Cody did tests with another VME (not VXS) crate and found the same problems with the fADC. They also found that when you put the fADC with an extension card at the last slot that doesn't work (#12), i.e. increasing the length of the VME bus, it worked better; so it's a matter of signal propagation in the bus. Next they will try to lower the fADC frequency from 80 to 40MHz.


- We are almost ready with the draft version of the TDR. Some numbers remain to be checked, Bill will help also with the main parameters of the cooling system and with one picture of the FDC cell. We discussed if the cooling system for all the detectors should be covered in a separate section. As for the FDC and CDC gas system, certainly it's worth having a separate section. Latest TDR version is linked above.